[Hallc_running] [New Comment] Hall C Status Update

latif at jlab.org latif at jlab.org
Mon Jan 22 20:35:02 EST 2018

 - by latif on Mon, 01/22/2018 - 20:34.

Hall C Status: Update on Target Replacement - The target group discovered that the carbon target clamp on the downstream side of the target melted (see attachment) and the carbon target dropped out of the ladder onto the base of the chamber. New target is being machined.
The target group is also making new target frames out of carbon instead of aluminum, which will hold the 6% carbon target and this will go into the scattering chamber tomorrow morning.  If time permits they will try to put in one additional spare thick carbon target replacing either the 0.5% thin carbon or the Halo target.


Logentry Text:
The target people went in and found the target frame has melted. They will be removing the frame.
Based on the current situation,  today's swing shift and owl shift  are cancelled. The day shift tomorrow will remain on standby. 


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3519135
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180122/9b075afa/attachment.html>
-------------- next part --------------
A non-text attachment was scrubbed...
Name: melted_target_frame.jpg
Type: image/jpeg
Size: 3568 bytes
Desc: not available
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180122/9b075afa/attachment.jpg>

More information about the Hallc_running mailing list